OpenStudy (anonymous):

Need someone to confirm??? Given f(x) = (x-1)/2, Solve for f^-1(5) a) 2 b) 9 c) 11 d) 15 I believe its C, can anyone help confirm?

OpenStudy (anonymous):

5 = (x-1)/2 solve for x

OpenStudy (shamil98):

f^-1 is x = y-1/2 2x + 1 = y 2(5) + 1 = y 11 = y so yes you are right.

OpenStudy (anonymous):

Thank you guys! ^^I might need some more help, but just with confirming answers if thats ok.